OpenStudy (anonymous):

I need help with math please wait while I type the equation into the comments.

OpenStudy (anonymous):

\[(-27)^{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

It's multiple choice a. -9 b. 9 c. 81 (I thought it was -9 but that turned out wrong)

OpenStudy (anonymous):

@AutumnRoseT thanks but I need to actually know how to do it and I don't need to graph it

OpenStudy (deana99):

I can help

OpenStudy (anonymous):

OK thank you @deana99 please do.

OpenStudy (anonymous):

Look, I know the answer is nine but I just need to know how to figure that out.

OpenStudy (amistre64):

the top of the fraction tells us the power, the bottom tell us the root

OpenStudy (amistre64):

what is (-27)^2

OpenStudy (anonymous):

729

OpenStudy (amistre64):

good now we root it \[\sqrt[3]{729}\]

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (amistre64):

2/3 power of 2, root of 3

OpenStudy (amistre64):

well, you can formulate a table of perfect cubes maybe ... and compare

OpenStudy (amistre64):

what is 10^3?

OpenStudy (anonymous):

300

OpenStudy (amistre64):

no, 10*10*10 is not 300

OpenStudy (here_to_help15):

hint 100*10= ?

OpenStudy (anonymous):

i mean 1000

OpenStudy (here_to_help15):

Ding ding :)

OpenStudy (amistre64):

1000 yes, and 729 is less than that, so we are less then 10^3 what is 5^3?

OpenStudy (anonymous):

125

OpenStudy (anonymous):

ok so it's nine

OpenStudy (anonymous):

because nine cubed is 729

OpenStudy (amistre64):

and 125 is less then 729 we know we are between 5 and 10 what is 6^3, 7^3, 8^3, and 9^3 yeah, its 9, since 9^3 = 729 :)

OpenStudy (anonymous):

OK thank you

OpenStudy (amistre64):

the idea tho is to power the top , then root the bottom

OpenStudy (amistre64):

one idea, if you feel spirited: (-27)^(2/3) = y (-27)^(2/3) * (-27)^(1/3) = y(-27)^(1/3) (-27)^(2/3+1/3) = y(-27)^(1/3) -27 = y(-3) -27/(-3) = y

OpenStudy (here_to_help15):

Dividing a negative and a negative you will get a positive! IMPORTANT TO KEEP IN MIND :)

OpenStudy (here_to_help15):

Positive and negative will be NEGATIVE AND + and + will be +

OpenStudy (anonymous):

Yeah thanks @Here_to_Help15 very helpful

OpenStudy (here_to_help15):

:)